OpenStudy (anonymous):

substitution method 6x+8y=44 -2x+y=11

OpenStudy (gw2011):

y=2x+11 6x+8(2x+11)=44 6x+16x+88=44 22x=-44 x=-2 Substitute this value into one of the equations to find y -2x+y=11 (-2)(-2)+y=11 4+y=11 y=7 Therefore, x=-2 and y=7

OpenStudy (anonymous):

gw2011 thank you could you assist me on this one please 3x+5y=-3 -7x+y=45

OpenStudy (gw2011):

y=7x+45 3x+5(7x+45)=-3 3x+35x+225=-3 38x=-228 x=-6 Substitute this value into one of the equations and you get: -7x+y=45 (-7)(-6)+y=45 42+y=45 y=3 Therefore, x=-6 and y=3

OpenStudy (anonymous):

gw2011 you did it again I really appreciate your time i'm not a fast learner at this type of math thanks for having sympathy on me

OpenStudy (gw2011):

You're welcome and I'm glad I could help.